A bullet is shot at a block that is initially at rest that is hanging from a cord. The bullet has a mass of 20 grams and and an initial speed of 650 m/s. If the block has a mass of 5kg and the bullet inside the block after the collision, what is the maximum height the block can swing after the collision?
Answer #1

here,

the mass of bullet , m = 20 g = 0.02 kg

the initial speed of bullet , u = 650 m/s

the mass of block , M = 5 kg

let the final speed of bullet and block after the collison be v

using conservation of momentum

m * u = ( m + M) * v

0.02 * 650 = ( 0.02 + 5) * v

solving for v

v = 2.59 m/s

the maximum height the block can swing , hm = v^2 / 2g

hm = 2.59^2 /(2*9.81) m

hm = 0.34 m

the maximum height of the block is 0.34 m